# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ B.4 Full Subsets Evaluation Results
Celeb- DF-v2 Gender $F_{MEO}$ 1.764 8.434 10.889 0.584 2.701 2.377 2.645 13.511 2.78 1.312 0.997 1.584
$F_{DP}$ 6.072 7.227 0.405 6.219 8.541 7.023 7 9.706 6.693 6.071 7.104 6.023
$F_{OAE}$ 1.578 2.063 5.663 1.092 2.149 0.976 0.884 8.053 0.636 2.411 0.831 0.429
$F_{EO}$ 2.585 10.238 11.073 1.108 3.236 3.379 3.564 20.484 3.369 1.599 1.519 1.601
Race $F_{MEO}$ 5.583 9.879 14.539 7.288 8.943 9.753 4.16 32.306 4.502 21.999 8.275 7.45
$F_{DP}$ 19.474 19.627 14.411 21.812 24.643 16.882 20.953 12.222 21.337 22.694 24.787 16.744
$F_{OAE}$ 6.569 9.664 12.618 4.032 6.035 5.493 3.524 10.82 3.813 3.092 5.392 2.815
$F_{EO}$ 10.691 25.652 28.759 13.013 11.726 15.42 14.684 63.524 9.671 58.225 14.714 12.384
Age $F_{MEO}$ 7.172 7.331 15.248 6.974 1.948 8.784 3.873 29.904 3.539 5.903 2.508 5.968
$F_{DP}$ 33.004 25.16 6.737 33.891 33.072 33.648 32.236 18.794 32.986 24.932 34.577 32.264
$F_{OAE}$ 1.925 8.576 26.359 1.628 1.532 1.149 2.502 12.526 3.482 10.577 0.845 1.183
$F_{EO}$ 11.497 14.073 19.966 11.657 5.013 11.178 7.669 53.72 9.685 10.027 5.404 7.037
Intersection $F_{MEO}$ 20 32.79 57.779 14.286 28.571 16.19 14.286 58.368 16.774 25.477 14.286 12.381
$F_{DP}$ 76.368 78.595 67.795 76.672 77.839 77.371 76.863 67.188 76.881 77.935 75.761 77.349
$F_{OAE}$ 19.231 16.228 49.562 11.538 8.463 7.334 7.692 29.689 11.538 5.769 7.692 5.769
$F_{EO}$ 71.129 114.538 103.126 53.655 59.887 59.694 60.765 182.729 61.653 141.381 48.621 33.495
- ACC 97.43 95.129 91.548 98.145 97.511 98.073 98.263 88.191 98.221 96.073 98.405 98.754
AUC 99.345 97.548 96.504 99.652 99.579 99.448 99.684 83.086 99.685 98.377 99.702 99.815
AP 99.908 99.641 99.492 99.953 99.943 99.923 99.957 97.068 99.957 99.763 99.96 99.974
EER 3.733 8.041 9.747 2.189 2.074 2.857 2.051 25.184 2.143 6.382 1.636 2.281
Table 19: Detailed fairness and utility evaluation results on Celeb-DF-v2.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
AttGAN Gender $F_{MEO}$ 0.56 1.669 0.472 0.946 0.459 1.554 0.422 4.79 3.544 1.941 0.153 1.47
$F_{DP}$ 10.923 12.249 10.678 11.069 11.068 10.489 11.295 9.998 12.045 11.936 11.171 12.205
$F_{OAE}$ 0.619 0.287 0.739 1.053 0.432 1.136 0.085 3.38 0.75 0.721 0.165 0.288
$F_{EO}$ 1.096 3.05 0.816 1.695 0.617 2.078 0.676 6.049 3.62 2.379 0.177 2.275
Race $F_{MEO}$ 3.39 3.613 3.228 3.198 3.918 4.013 3.03 18.643 17.655 7.576 1.887 1.695
$F_{DP}$ 11.859 13.88 11.05 12.876 10.628 11.582 13.059 22.615 16.834 13.753 13.054 13.502
$F_{OAE}$ 1.587 2.174 2.526 2.387 2.033 2.31 2.521 4.713 5.636 5.042 1.6 1.6
$F_{EO}$ 5.016 10.539 9.994 5.975 9.472 9.239 5.592 37.89 19.269 12.917 4.291 5.117
Age $F_{MEO}$ 3.086 4.899 7.144 1.194 3.704 3.096 2.469 15.211 5.996 2.855 2.206 5.493
$F_{DP}$ 22.439 21.386 18.175 23.439 22.14 22.789 24.491 20.473 21.175 22.14 24.789 21.193
$F_{OAE}$ 1.105 3.595 4.689 0.942 2.563 3.132 2.456 6.436 4.493 2.309 0.398 3.758
$F_{EO}$ 5.209 10.255 13.136 4.103 10.371 10.312 5.807 36.092 8.639 6.932 3.746 7.407
Intersection $F_{MEO}$ 5.128 11.111 11.111 7.692 7.407 6.667 7.407 31.774 33.333 7.692 3.125 5
$F_{DP}$ 20.594 24.253 20.152 21.106 20.783 19.375 22.003 28.514 28.753 21.677 22.003 23.411
$F_{OAE}$ 4.225 7.042 4.968 5.634 3.177 4.878 4.348 16.17 10.976 5.479 2.817 1.852
$F_{EO}$ 21.471 42.762 35.107 23.6 25.943 31.389 22.546 92.264 46.215 33.053 12.897 17.18
- ACC 98.482 97.884 95.86 98.62 98.482 98.666 98.712 80.957 96.274 97.608 99.264 99.126
AUC 99.798 99.526 99.259 99.776 99.702 99.642 99.875 89.719 98.721 99.722 99.781 99.953
AP 99.795 99.492 99.282 99.797 99.612 99.587 99.888 91.76 98.646 99.732 99.827 99.958
EER 1.594 2.092 4.084 1.494 1.494 1.394 1.195 18.426 4.98 2.39 0.996 1.494
Table 20: Detailed fairness and utility evaluation results on AttGAN.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
MMDGAN Gender $F_{MEO}$ 2.722 17.144 0.773 1.087 3.809 3.261 4.622 1.626 3.394 8.439 11.417 2.448
$F_{DP}$ 8.077 16.007 6.925 7.801 8.077 7.939 9.604 6.787 9.584 11.082 13.928 9.141
$F_{OAE}$ 2.335 7.348 1.084 1.271 3.398 3.26 2.797 0.63 0.512 4.275 4.994 1.133
$F_{EO}$ 4.772 18.286 0.773 2.028 6.9 6.352 5.63 3.071 4.481 9.447 12.493 2.448
Race $F_{MEO}$ 10 73.95 8.974 16.667 14.286 8.333 8.333 33.333 16.667 28.571 28.571 2.21
$F_{DP}$ 33 55 32 33 38 33 33 33 33 43 43 33
$F_{OAE}$ 9.091 22 5.233 9.091 5 4.545 4.545 18.182 9.091 10 10 1.187
$F_{EO}$ 23.462 91.198 13.141 22.352 28.661 15.508 19.122 48.478 25.201 39.585 44.593 5.931
Age $F_{MEO}$ 11.706 22.297 4.808 10.345 10.345 10.345 7.642 6.924 9.091 14.336 14.559 9.091
$F_{DP}$ 17.703 20.303 10.909 9.394 11.515 11.212 11.818 10.303 12.727 8.788 12.727 11.818
$F_{OAE}$ 3.939 10.606 2.424 5.455 9.091 6.515 4.127 3.828 2.233 6.89 8.254 5.263
$F_{EO}$ 23.368 39.483 6.422 13.465 22.203 25.61 17.407 13.78 17.996 21.652 24.926 10.142
Intersection $F_{MEO}$ 12.5 100 22.222 22.222 16.667 11.111 11.111 44.444 22.222 100 33.333 4.167
$F_{DP}$ 58.088 62.5 51.471 58.088 58.088 58.088 58.088 58.088 58.088 70.588 58.088 58.088
$F_{OAE}$ 11.765 41.667 12.5 11.765 8.333 5.882 5.882 23.529 11.765 12.5 16.667 1.948
$F_{EO}$ 51.536 230.67 59.507 44.743 59.982 39.11 42.671 103.142 55.926 161.146 85.72 14.412
- ACC 97.525 90.099 95.792 98.02 97.03 98.02 97.772 93.812 96.535 94.307 96.287 99.01
AUC 99.395 97.839 99.299 99.687 99.508 99.392 99.781 97.987 98.521 99.515 99.808 99.98
AP 98.918 97.589 99.226 99.691 99.461 99.215 99.792 97.182 98.25 99.525 99.812 99.983
EER 2.646 7.407 3.704 1.587 2.646 2.646 2.116 6.349 4.233 1.587 1.587 0.529
Table 21: Detailed fairness and utility evaluation results on MMDGAN.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
StarGAN Gender $F_{MEO}$ 0.281 2.181 0.472 0.33 0.388 0.552 0.044 2.075 1.737 0.344 0.627 0.33
$F_{DP}$ 4.379 5.639 3.887 4.547 4.555 4.744 4.461 5.32 5.209 4.678 4.954 4.65
$F_{OAE}$ 0.181 0.604 0.435 0.3 0.275 0.332 0.008 0.095 0.214 0.07 0.107 0.197
$F_{EO}$ 0.408 2.679 0.59 0.486 0.561 0.619 0.049 2.74 2.18 0.606 0.957 0.375
Race $F_{MEO}$ 4 4.577 11.031 4 8 4 8 22.727 11.197 6.113 4 2.062
$F_{DP}$ 27.875 27.493 29.459 28.39 29.39 26.875 29.086 35.768 30.974 27.403 27.875 26.056
$F_{OAE}$ 1.515 3.036 4.571 1.515 3.03 3.03 3.03 6.682 2.931 1.395 1.515 1.325
$F_{EO}$ 7.666 11.827 17.243 8.507 10.863 8.762 10.006 29.26 17.857 12.09 6.291 4.109
Age $F_{MEO}$ 2.479 3.577 5.091 3.167 1.667 2.5 1.379 1.167 4.562 2.033 1.667 0.943
$F_{DP}$ 19.078 17.476 16.434 19.802 19.399 19.078 19.319 17.244 18.927 17.39 19.078 19.158
$F_{OAE}$ 1.132 2.264 2.119 0.323 1.201 2.264 1.132 2.075 0.843 0.908 1.509 0.601
$F_{EO}$ 4.659 6.124 9.058 5.728 4.038 5.801 2.946 2.421 6.082 4.539 3.86 2.745
Intersection $F_{MEO}$ 14.286 12.5 21.774 6.25 11.111 6.25 14.286 25 18.75 11.111 5.556 2.381
$F_{DP}$ 30.971 32.154 36.599 31.973 33.612 29.932 31.571 38.639 36.417 31.791 31.571 28.326
$F_{OAE}$ 5.882 5.882 7.418 2.222 4.082 4.082 5.882 8.889 5.462 4.082 2.041 1.471
$F_{EO}$ 22.432 38.089 41.688 20.567 19.704 19.605 25.484 59.844 36.081 26.652 11.756 8.426
- ACC 99.326 98.289 96.216 99.015 99.274 99.378 99.43 94.66 96.319 98.237 99.482 99.533
AUC 99.874 99.773 99.556 99.909 99.86 99.869 99.964 99.626 99.076 99.796 99.909 99.983
AP 99.899 99.797 99.56 99.933 99.826 99.832 99.97 99.724 99.079 99.809 99.929 99.986
EER 0.795 1.135 2.611 0.454 0.795 0.681 0.568 1.93 3.973 1.589 0.568 0.454

### v3

@@ D.1 Detailed Results of Overall Performance Comparison
Fairness(%) Skin Tone $F_{MEO}$ 15.463 6.826 7.442 13.642 4.221 9.425 13.574 2.368 13.487 10.127 8.763 7.613
$F_{DP}$ 11.994 10.440 8.378 11.211 8.998 6.875 11.085 8.231 11.506 10.182 8.920 11.914
$F_{OAE}$ 1.998 2.390 1.540 1.713 2.141 7.511 1.661 2.777 1.538 1.375 1.484 1.822
$F_{EO}$ 0.192 0.107 0.095 0.171 0.064 0.187 0.168 0.036 0.170 0.119 0.114 0.129
Gender $F_{MEO}$ 4.209 3.639 9.461 5.189 4.116 2.328 5.402 4.084 5.058 4.112 4.143 4.035
$F_{DP}$ 1.171 1.043 3.025 1.749 1.191 2.803 1.778 1.082 1.899 1.537 1.560 1.401
$F_{OAE}$ 1.579 1.395 2.163 1.499 1.507 1.248 1.560 1.519 1.277 1.218 1.158 1.353
$F_{EO}$ 0.051 0.045 0.095 0.057 0.050 0.037 0.059 0.050 0.053 0.045 0.045 0.046
Age $F_{MEO}$ 33.930 16.272 10.222 45.076 20.048 11.857 45.508 5.788 37.055 30.409 29.707 20.058
$F_{DP}$ 15.167 11.254 11.643 15.938 11.379 8.360 15.357 15.620 13.565 11.925 9.880 10.487
$F_{OAE}$ 8.520 4.738 5.662 11.409 5.123 10.447 11.546 3.777 9.454 7.784 8.513 4.450
$F_{EO}$ 0.488 0.286 0.231 0.623 0.322 0.228 0.625 0.136 0.526 0.441 0.459 0.357
Intersection $F_{MEO}$ 19.488 11.396 16.807 15.829 6.631 12.599 16.926 5.116 16.594 13.597 11.226 10.523
$F_{DP}$ 17.093 15.841 11.946 15.708 13.376 11.010 15.607 13.121 15.834 14.758 12.474 16.906
$F_{OAE}$ 3.583 5.700 5.469 3.139 5.452 10.020 2.878 6.209 2.852 2.313 2.755 4.348
$F_{EO}$ 0.392 0.262 0.340 0.349 0.225 0.420 0.342 0.188 0.345 0.254 0.226 0.294
Utility(%) - AUC 0.985 0.985 0.987 0.981 0.984 0.979 0.982 0.981 0.984 0.981 0.987 0.988
ACC 0.950 0.949 0.940 0.958 0.950 0.816 0.956 0.936 0.959 0.963 0.960 0.953
AP 0.994 0.994 0.996 0.992 0.994 0.992 0.993 0.994 0.993 0.989 0.995 0.996
EER 0.065 0.064 0.062 0.069 0.064 0.066 0.068 0.078 0.060 0.053 0.057 0.058
FPR 0.145 0.162 0.206 0.134 0.148 0.039 0.141 0.189 0.138 0.118 0.144 0.099
Table D.5: Detailed fairness and utility evaluation results on 80% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 9.750 11.908 5.095 12.896 5.859 18.083 10.363 5.926 10.955 9.570 8.980 3.889
$F_{DP}$ 16.824 14.637 12.547 17.621 14.479 19.676 15.803 8.744 16.109 16.063 15.700 12.899
$F_{OAE}$ 2.401 2.973 1.714 3.630 0.905 5.752 3.249 5.302 3.524 2.641 2.684 1.164
$F_{EO}$ 0.164 0.155 0.119 0.183 0.113 0.248 0.147 0.083 0.149 0.142 0.132 0.069
Gender $F_{MEO}$ 3.535 0.952 4.275 4.788 4.439 4.201 4.655 6.898 3.981 3.848 3.674 3.792
$F_{DP}$ 2.241 0.715 3.129 3.184 3.060 2.245 3.426 4.328 2.850 2.562 2.514 2.790
$F_{OAE}$ 2.482 0.833 2.263 2.690 2.496 2.920 2.332 3.028 2.236 2.451 2.360 2.166
$F_{EO}$ 0.051 0.011 0.047 0.057 0.052 0.062 0.049 0.070 0.048 0.051 0.049 0.045
Age $F_{MEO}$ 23.282 14.195 13.627 32.679 26.796 52.939 30.332 8.251 32.612 23.934 23.834 15.408
$F_{DP}$ 17.521 18.304 16.426 20.437 18.489 28.303 18.892 12.343 19.913 17.103 18.221 15.255
$F_{OAE}$ 11.063 3.474 9.185 16.411 13.102 26.805 15.234 6.602 16.794 11.909 11.523 7.893
$F_{EO}$ 0.381 0.203 0.303 0.490 0.436 0.769 0.472 0.162 0.469 0.378 0.370 0.273
Intersection $F_{MEO}$ 11.921 17.564 7.727 17.622 10.201 21.679 15.031 12.980 12.775 13.010 11.104 7.173
$F_{DP}$ 23.484 21.311 17.501 23.519 21.115 24.119 22.167 13.434 21.787 22.542 22.120 18.247
$F_{OAE}$ 4.000 4.650 3.969 5.957 4.306 9.974 5.003 11.530 5.750 3.903 4.600 4.426
$F_{EO}$ 0.344 0.335 0.304 0.378 0.250 0.515 0.324 0.270 0.325 0.290 0.303 0.196
Utility(%) - AUC 0.984 0.984 0.984 0.982 0.982 0.935 0.981 0.979 0.986 0.983 0.987 0.989
ACC 0.937 0.892 0.922 0.928 0.929 0.900 0.930 0.830 0.933 0.937 0.944 0.940
AP 0.980 0.986 0.986 0.977 0.978 0.910 0.977 0.980 0.982 0.980 0.984 0.987
EER 0.062 0.064 0.067 0.061 0.066 0.129 0.065 0.082 0.056 0.061 0.053 0.052
FPR 0.087 0.005 0.111 0.122 0.111 0.164 0.116 0.337 0.115 0.095 0.083 0.095
Table D.6: Detailed fairness and utility evaluation results on a training subset with the ratio of real vs fake is 1:1.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 11.678 10.565 8.595 12.629 11.790 17.068 9.661 5.615 11.138 11.726 8.680 6.435
$F_{DP}$ 14.133 12.859 10.579 15.157 13.983 16.962 14.081 8.438 14.161 14.724 13.388 13.256
$F_{OAE}$ 4.539 3.671 4.407 4.104 4.894 5.036 3.931 5.461 4.389 3.379 3.006 2.232
$F_{EO}$ 0.128 0.151 0.102 0.148 0.127 0.200 0.107 0.081 0.121 0.141 0.103 0.081
Gender $F_{MEO}$ 6.942 2.295 10.586 7.818 7.327 8.990 6.054 6.525 7.518 6.259 5.934 4.944
$F_{DP}$ 4.378 0.094 6.203 4.881 4.632 5.572 4.086 4.136 4.867 4.093 3.917 3.565
$F_{OAE}$ 3.225 1.508 4.799 3.669 3.303 4.060 2.906 2.842 3.445 2.890 2.795 2.407
$F_{EO}$ 0.072 0.023 0.106 0.080 0.074 0.091 0.063 0.067 0.075 0.065 0.062 0.052
Age $F_{MEO}$ 36.384 13.175 25.574 35.508 30.942 36.806 32.134 6.629 34.717 32.860 29.474 28.923
$F_{DP}$ 18.815 19.006 16.522 18.393 17.007 18.947 17.086 12.524 18.454 16.798 15.004 19.634
$F_{OAE}$ 19.144 2.128 14.899 18.668 16.331 19.294 16.714 6.598 18.249 17.426 15.605 15.099
$F_{EO}$ 0.524 0.193 0.373 0.525 0.442 0.553 0.490 0.147 0.507 0.482 0.447 0.420
Intersection $F_{MEO}$ 16.037 18.196 19.081 16.394 19.895 21.535 13.589 12.135 15.921 14.772 12.424 10.340
$F_{DP}$ 16.749 17.705 16.207 17.525 18.313 19.312 17.029 12.639 17.144 16.813 15.850 16.846
$F_{OAE}$ 7.914 4.565 12.301 6.936 8.704 9.372 5.877 11.469 7.025 5.523 6.909 5.109
$F_{EO}$ 0.381 0.346 0.400 0.399 0.394 0.467 0.317 0.249 0.364 0.336 0.313 0.243
Utility(%) - AUC 0.958 0.967 0.975 0.966 0.964 0.951 0.976 0.978 0.969 0.962 0.967 0.983
ACC 0.864 0.864 0.823 0.876 0.855 0.862 0.909 0.829 0.888 0.874 0.882 0.925
AP 0.934 0.972 0.976 0.948 0.946 0.938 0.964 0.979 0.952 0.942 0.951 0.977
EER 0.087 0.098 0.091 0.082 0.081 0.116 0.064 0.085 0.073 0.084 0.079 0.056
FPR 0.267 0.012 0.349 0.242 0.285 0.272 0.172 0.339 0.221 0.246 0.227 0.142
Table D.7: Detailed fairness and utility evaluation results on a training subset with the ratio of real vs fake is 1:10.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 6.557 11.357 7.743 7.364 5.850 18.290 5.814 12.493 6.397 5.083 6.515 5.122
$F_{DP}$ 13.087 13.112 12.348 15.751 11.728 22.766 14.881 13.771 14.364 12.265 15.536 12.718
$F_{OAE}$ 1.779 2.939 1.589 1.352 2.402 3.823 1.495 2.386 0.996 1.401 1.471 1.742
$F_{EO}$ 0.108 0.140 0.093 0.153 0.090 0.331 0.135 0.155 0.125 0.084 0.143 0.096
Gender $F_{MEO}$ 1.808 1.803 2.106 1.659 2.316 3.035 3.035 1.997 3.499 1.920 1.812 3.016
$F_{DP}$ 1.769 0.136 0.369 1.321 0.470 2.871 2.503 2.008 2.119 1.097 1.405 2.633
$F_{OAE}$ 1.385 1.269 1.615 1.661 1.777 1.270 1.817 0.579 2.662 1.928 1.734 1.677
$F_{EO}$ 0.025 0.018 0.027 0.032 0.032 0.038 0.035 0.020 0.052 0.036 0.033 0.033
Age $F_{MEO}$ 8.571 11.954 9.832 8.680 9.523 36.509 11.812 10.258 13.180 8.539 9.553 11.109
$F_{DP}$ 17.740 18.338 18.662 17.788 18.910 27.851 16.998 16.446 16.719 18.069 18.034 16.696
$F_{OAE}$ 3.191 0.656 1.405 4.291 2.010 12.139 5.174 2.737 5.723 3.924 4.283 4.859
$F_{EO}$ 0.196 0.157 0.141 0.228 0.174 0.685 0.281 0.173 0.288 0.209 0.240 0.246
Intersection $F_{MEO}$ 11.785 18.146 14.840 14.116 12.940 23.975 11.313 16.240 12.729 10.948 13.399 9.441
$F_{DP}$ 20.192 17.674 19.309 23.135 18.594 29.625 22.394 19.583 21.935 18.674 23.584 17.906
$F_{OAE}$ 3.419 4.233 3.638 4.263 4.020 5.166 3.479 3.133 4.945 3.388 3.885 3.667
$F_{EO}$ 0.251 0.312 0.256 0.329 0.232 0.678 0.305 0.307 0.285 0.225 0.314 0.232
Utility(%) - AUC 0.978 0.973 0.982 0.979 0.982 0.933 0.978 0.975 0.979 0.980 0.982 0.983
ACC 0.920 0.862 0.895 0.928 0.916 0.832 0.921 0.849 0.921 0.920 0.930 0.933
AP 0.978 0.977 0.984 0.978 0.984 0.915 0.979 0.978 0.978 0.979 0.981 0.984
EER 0.070 0.088 0.075 0.066 0.064 0.141 0.076 0.087 0.074 0.070 0.065 0.066
FPR 0.034 0.008 0.009 0.042 0.018 0.116 0.054 0.004 0.055 0.037 0.040 0.051